1. Consider the reaction: 5 A + 3 B → 2 C The rate of the reaction is found to be 0.0486 M/s. What is the rate of change of B? Be sure to consider whether the substance is disappearing or appearing in your answer.

2. For the following zero order reaction, the rate constant is 0.0311. If [A]0 = 8.68, calculate [A] at 2.9 seconds.

2 A →→ 3 B
Answer #1

1)

rate of reaction is given as:

rate = sign * (1/x) * d[concentration]/dt

sign is negative for reactant and positive for product

x is stoichiometric coefficient in chemical equation

so, rate expression in terms of B is:

Rate = (-1/3)*Δ[B]/Δt

0.0486 M/s = (-1/3)*Δ[B]/Δt

Δ[B]/Δt = -0.146 M/s

Answer: -0.146 M/s

2)

we have:

[A]o = 8.68 M

t = 2.9 s

k = 3.11*10^-2 M/s

Given:

[A]o = 8.68 M

use integrated rate law for 0 order reaction

[A] = [A]o - k*t

[A] = 8.68 - 3.11*10^-2*2.9

= 8.59 M

Answer: 8.59 M